You wish to cross a River from point A to the other side of the river, point B. You are using a boat that can travel at 7mph, the river has a current of 2mph. It’ll take you 15 minutes to cross the river, how far downstream will you be by the time you reach point B?
I hope I’m doing this right. I got 1.75(2)=3.5, but that answer isn’t an answer choice.
Thank you!

Found 2 solutions by math_helper, ikleyn:
Answer by math_helper(2461)


Let's say the boat is heading from West to East. It is given that the boat will spend 15mins (=1/4hr) crossing. If the boat is continuously pointed directly East, then the only additional information you need is the 2mi/hr river current.
Distance downstream upon landing at East side of river is:
2mi/hr * (1/4)hr = (1/2)mi

Answer by ikleyn(52788)

.

In this post,  two different problems are mixed - - - which makes the whole " problem "  INVALID.

One problem is about crossing the river using a boat,  from one its bank to another.

The other problem is floating downstream at given boat's own speed and added current speed.


When two different problems are mistakenly mixed in one,
the resulting product is only good   to  WHAT ?     to  WHAT ? ?     to  WHAT ? ? ?


        OK,  in this post my task is to disguise this defective problem's formulation
        (for future generation of students,  who will read it - to save their time and their brain . . . )


It is a surprise to me that the other tutor does not see it . . .